OpenStudy (lukecrayonz):

Write each degree measure in radians as a multiple of pi and each radian measure in degrees. 1. 120 degrees 2. 5pi/6

OpenStudy (anonymous):

1. 2 pi / 3 2. 150

OpenStudy (lukecrayonz):

how :O

OpenStudy (anonymous):

\[120 \times \frac{\pi}{180} = \frac{2 \pi } {3}\]

OpenStudy (lukecrayonz):

I know you do pi/180 and 180/pi, but not sure for which

OpenStudy (lukecrayonz):

so can i do 120/180, then put it into my calculator, and use the fraction button, get 2/3, then just put pi on top?

OpenStudy (anonymous):

yes

OpenStudy (lukecrayonz):

and so for the 5pi/6, i do 180/pi 5pi/6 * 180/pi

OpenStudy (lukecrayonz):

\[5\pi/6*180/\pi\]

OpenStudy (anonymous):

yes so the pi cancels out

OpenStudy (lukecrayonz):

i can just put that into the calculator and get 150, right?